OpenStudy (anonymous):

How do I find the solution of x^2 - 81 using square roots? I need help >.<

OpenStudy (luigi0210):

Start by getting \(x^2\) by itself

OpenStudy (anonymous):

Isn't it by itself already?

OpenStudy (luigi0210):

Nope, there's a -81 next to it o:

OpenStudy (anonymous):

\[-81x^2\]?

OpenStudy (anonymous):

@Luigi0210

OpenStudy (luigi0210):

No.. like this. \[\LARGE x^2-81=0\] Get x^2 by itself \[\LARGE x^2=81\] Now take the sqaure root of both sides, what do you get?

OpenStudy (anonymous):

The square root of 81 is 9

OpenStudy (anonymous):

I'm not sure what the square root of \[x^2\]

OpenStudy (anonymous):

\[\sqrt{x ^{2}} \] is just x because they cancels out

OpenStudy (anonymous):

What do I do next?

OpenStudy (anonymous):

you already got the answer x=9

OpenStudy (anonymous):

Well... My answer choices are -9(squared) and 9(squared) -81 and 81 -9 and 9 no real number solutions

OpenStudy (anonymous):

Is it option C.?

OpenStudy (anonymous):

Not squared sorry, I meant square rooted in the parenthesis.

OpenStudy (anonymous):

yes if you meant -9 and 9

OpenStudy (anonymous):

Okay thanks :)

OpenStudy (anonymous):

as \[\sqrt{81} = \pm9\]

OpenStudy (anonymous):

Could you help me with one more?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

\[3x^2-147\]

OpenStudy (anonymous):

Isolate the x?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[x^2=147(3)\]?

OpenStudy (anonymous):

uhh nope

OpenStudy (anonymous):

you isolate by subtracting 147 first\[3x ^{2} =147\]

OpenStudy (anonymous):

Subtracting 147-3?

OpenStudy (anonymous):

to bring the 3 to the other side, you do the opposite what is the opposite of multiplying by 3?

OpenStudy (anonymous):

Dividing, okay so 147/3=49

OpenStudy (anonymous):

Then square root both so the square root of 49 is 7

OpenStudy (anonymous):

Ohhhh it's easy :D I'm just over complicating things.

OpenStudy (anonymous):

yes! but careful \[\sqrt{49} = \pm 7 \]

OpenStudy (anonymous):

What do you mean be careful?

OpenStudy (anonymous):

usually a question asking for sqrt of 49 would be 7, but then it could also be -7 at the same time as -7x-7 also equals 49

OpenStudy (anonymous):

Yeah but I'm okay because the answer choice is -7 and 7.

OpenStudy (anonymous):

yes correct

OpenStudy (anonymous):

Yayy, thank you for your help!

OpenStudy (anonymous):

no problem